# The Language of Interoception: Examining Embodiment and Emotion Through a Corpus of Body Part Mentions

### v1

@@ Abstract
This paper is the first investigation of the connection between emotion, embodiment, and everyday language in a large sample of natural language data. We created corpora of body part mentions (BPMs) in online English text (blog posts and tweets). This includes a subset featuring human annotations for the emotions of the person whose body part is mentioned in the text. We show that BPMs are common in personal narratives and tweets ($\sim$5to 10% of posts include BPMs) and that their usage patterns vary markedly by time and location. Using word–emotion association lexicons and our annotated data, we show that text containing BPMs tends to be more emotionally charged, even when the BPM is not explicitly used to describe a physical reaction to the emotion in the text. Finally, we discover a strong and statistically significant correlation between body-related language and a variety of poorer health outcomes. In sum, we argue that investigating the role of body-part related words in language can open up valuable avenues of future research at the intersection of NLP, the affective sciences, and the study of human wellbeing.
@@ 1 Introduction
Embodied cognition—the theory that human cognition is rooted in bodily experiences—has gained significant traction across cognitive science, psychology, linguistics, and philosophy. This framework suggests that our bodily experiences shape not only how we interact with our physical environment, but also how we process, represent, and share abstract concepts. A wide range of disciplines have demonstrated that the intersection of cognition and language is deeply intertwined with sensorimotor experiences. For example, a growing pool of research suggests that individuals learn new concepts better when they can use their bodies to simulate the concepts Cook et al. (2008); Johnson-Glenberg et al. (2016), language processing involves mental simulation of physical actions Pulvermüller (2005); Glenberg and Kaschak (2002), and emotions emerge from interpretations of physiological signals through a process known as interoception Craig (2002); Barrett (2017).
While it is well established that embodiment holds a key role in human cognition, an important open question is the extent to which our embodied experiences are encoded in and reflected through everyday language. NLP corpora and techniques offer a unique opportunity to systematically investigate this relationship at scale. To this end, we introduce the concept of Body Part Mentions (BPMs), which we define as instances of language where words associated with parts of the body are used. While BPMs are founded on a relatively simple connection between language and embodiment—words with a semantic association to the human body—we argue that they offer a new lens with which we can investigate the deep relationship between language and our embodied selves.
A better understanding of embodiment through language could support a range of research areas, including: NLP in health domains (where BPM-heavy corpora are frequent and their interpretations carry severe consequences for patients Chaturvedi et al. (2023)), the growing wave of interest in integrating embodiment within computational models of language (especially text-based LLMs, which rely on BPMs to access information about human embodiment, and struggle with many benchmarks of human cognition due to their lack of embodiment Chemero (2023)), and cross-linguistic and cross-cultural studies on embodiment (which are often limited by difficulties in accessing a wide range of participants from different backgrounds) Kóczy and Sipőcz (2023).
@@
In this work, we introduce two novel BPM corpora, Spinn3rBPM (a corpus of blog posts) and TUSCBPM (a corpus of tweets), and conduct a variety of experiments to show how they can be used to address three fundamental questions about the connection between embodiment and natural language. First, we investigate the prevalence of body-part related words in everyday language and how different BPM categories differ in terms of the context that surrounds them. Second, we look at the relationship between BPMs and affect, motivated by the growing pool of research which posits that emotions originate from interpretations of physiological signals. Lastly, we propose and test the hypothesis that the extent of BPMs in social media is indicative of aggregate-level health outcomes.
@@ 2.1 Embodiment, Affect, and Health
The theory of constructed emotion, (introduced by Lisa Feldman Barrett Barrett (2017), a leading scholar in the affective sciences), proposes that emotions emerge as interpretations of our bodies’ physiological signals. A key argument for this theory is the wide range of research establishing a strong connection between interoception–the ability to feel internal bodily sensations Craig (2002)–and emotional wellbeing. Better interoceptive awareness has been shown to positively correlate with better emotional regulation Zamariola et al. (2019), emotional decision-making Dunn et al. (2010), and emotional granularity (the ability to distinguish different emotions) Feldman et al. (2024). These results indicate that awareness of our bodily experiences is crucial to our emotional welfare and dysfunctional interoception is possibly a major contributor to a variety of mental health conditions Khalsa et al. (2018).
Embodied experiences also manifest frequently in everyday language. For example, descriptions of bodily experiences are frequently found in storytelling contexts Gallese and Wojciehowski (2011). But even outside of their role as explicitly physical referents, body-related words may reveal a deeper connection to embodied phenomena. The theory of conceptual metaphor suggests that metaphors are a fundamental cognitive process Lakoff and Johnson (2008). According to this framework, metaphors help us understand abstract concepts by mapping them onto concrete, physical experiences. Treated this way, the widespread usage of body part words in language—even those applied in abstract contexts—are important for understanding how natural language is connected to the physical world from which it originates. Scholars in affective theory have argued that this is why intensely emotional experiences (e.g., “my heart is broken”, “weight lifted off my shoulders”) often use body parts and actions as a metaphor; they are effective ways of grounding subjective experience in a shared embodied reality Kövecses (2003).
@@ 2.2 Embodiment and NLP
Language plays a key role in understanding the relationship between embodiment and affect. Emotional granularity is measured from the ability to identify and distinguish different emotions using words Tan et al. (2022), and interoception is often measured through an individual’s ability to describe their internal state Desmedt et al. (2022). Notably, a range of NLP projects have taken an interest in body part words for specific applications, such as identifying gendered representations in literature Silva et al. (2024), mapping bodily sensations for healthcare applications Wang et al. (2019), or building computational methods for detecting body parts involved in emotional processes to improve machine emotion recognition Zhuang et al. (2024). Yet we did not find prior work with the goal of analyzing body-part related words in everyday language, or investigating whether phenomena that have been theorized and observed in laboratory environments (i.e., the relationship between affect and body part words suggested by the theory of conceptual metaphor, or the relationship between affect and described bodily experiences suggested by the theory of constructed emotion), can be corroborated using textual corpora.
In this paper, we primarily use lexicons for word-emotion association—which can be used to effectively distinguish emotional granularity Vishnubhotla et al. (2024) or create accurate emotion arcs in streams of text Teodorescu and Mohammad (2023)—as a computationally inexpensive and interpretable method for beginning to look at the relationship between affect, language, and embodiment.
@@ 3 BPM Corpora
For this work, we investigate two mediums for online, everyday language: blog posts and tweets. We consider each sentence from a blog post and each individual tweet from the tweet corpora as instances. We created BPM corpora by extracting all instances that included at least one word referring to a body part from the Spinn3r personal blog datasets Burton et al. (2009, 2011), and the two TUSC datasets: TUSCctry (where tweets are geo-located to either the United States of America or Canada) and TUSCcity (where tweets are geo-located to cities in North America) Vishnubhotla and Mohammad (2022). This results in three final BPM corpora: Spinn3rBPM, TUSCctry-BPM, and TUSCcity-BPM. We will refer to the corpora made up of the rest of the instances as Spinn3rnoBPM and TUSCnoBPM.
We use the same body parts word lists as in Zhuang et al. (2024), which extracts BPM samples from the Spinn3r corpus to annotate samples for the presence of explicitly embodied emotion. [^1] https://www.collinsdictionary.com/us/word-lists/body-parts-of-the-body [^2] https: //www.enchantedlearning.com/wordlist/body.shtmlAdditionally, we include plural forms of terms (e.g., hearts, hands, eyes, etc.) in our final BPM list. We refer to the body part word forms as BP word types. The list of 295 BP word types we used is available on the project webpage.
Two issues need to be addressed when working with BPM corpora to study embodiment. First, some words in the BPM word types list are ambiguous (some BPM instances may not actually be referring to a body part: e.g, ‘I will be back’). Second, it is useful to distinguish between the speaker referring to their own body vs. the speaker referring to someone else’s body—we are especially interested in mentions of one’s own body as a possibly useful indicator for well-being.
We use a simple solution that effectively addresses both of the issues raised above. We created three subsets for each BPM corpus that only include the BPM instances preceded by possessive pronouns ‘my’, ‘your’, and ‘his/her/their’, respectively, and call these instances: possessed BPMs This helps us create separate corpora for first-, second-, and third-person references to body parts, and also excludes a vast majority of mentions that are not explicit references to body parts. For example, ‘I will be back’ would not be included in any of these corpora, but ‘my back hurts’ will be. We find that this approach delivers a high number of references to an individual’s body part (92% of 100 manually inspected instances). While this approach sacrifices recall by excluding possible references to body parts without possessive pronouns, we benefit from higher precision. Additionally, we also conduct some experiments with all BPM instances (the higher-recall and lower-precision corpus).
@@ 3.1 Emotion-Annotated BPM Corpus
We also leverage a specialized subset of Spinn3r, previously released by Zhuang et al.. This subset contains sentences with BPMs annotated for the presence of explicitly embodied emotion, which is defined as "the physical experience of an emotion via our body" (i.e., "Julie pouted and rolled her eyes" is annotated as containing embodied emotion, but "Frank breathed heavily through his mouth after his run" is not). We refer to this subset in the rest of this paper as Spinn3rBPM-Zhuang. We extend this work by creating the first human-annotated dataset that explicitly identifies BPM ownership (whether the BPM refers to the speaker’s body or not) and the emotion of the BPM owner (joy, fear, etc.) as inferred by a human reader.
For full details of all datasets and data collection methods see Appendix A. Further details for the emotion annotation process, which includes a set of quality checks and aggregation methods for final emotion scores, can be found in Appendix B. All of our data is made freely available for research.
@@ 4 Research Questions About BPMs
Despite substantial evidence from medical research and psychology that points to connections between the mind and rest of the body, as well as the connection between interoception and emotional well-being, there is little quantitative work using language to explore this connection. In this section and the next, we make use of large amounts of social media data, massive word–emotion lexicons, and the emotion-annotated corpus described in the previous section, to examine questions on how, when, and in what context we refer to our body parts in text (this section) and whether different body parts mentions tend to be used in different emotional contexts (next section). Since the questions in this section are relevant to the Body, we will index them as B1, B2, etc. The questions in the next section are related to Body and Affect, so we will index them as BA1, BA2, etc.
@@
Method: We did not find any past research on how common BPMs are in language. We do not even have a sense of the magnitude: do they occur in 0.01%, 0.0001%, 1%, or 10% of utterances? To examine the extent to which body-related vocabulary is used in online language, we calculated the number of instances containing at least one BPM.
Results: See Table 1. We find that a substantial proportion of instances contain at least one BPM: 10.41% of blog post sentences in the Spinn3r dataset, 6.42% and 7.28% of tweets in in TUSCctry and TUSCcity, respectively.
Discussion: The consistently high proportion of BPM-containing instances across our corpora emphasizes the ubiquity of body part references in online English text. The markedly larger proportion samples containing BPMs in blog sentences than tweets implies that usage of BPMs may be more pervasive in personal narratives and longer-form text than in tweets.
@@
Method: Although BPMs can be used to study the general significance of body part words in language, we are also interested in examining the extent to which these body parts can be attributed to a particular person (e.g., her heart) vs. body part words that are not attributed to a human possessor (e.g., the heart of the matter). To do this, we introduce the concept of possessed BPMs, which can be attributed to a particular body using a possessive pronoun. We look at three general categories of possessed BPMs: first person instances including "my <BPM>", second person instances including "your <BPM>", and third person instances including "his/her/their <BPM>". In each corpus, we determine the frequency of instances containing at least one instance of a possessed BPM.
Results: Table 2 shows the results. In each corpus, the most common possessed BPM is my <BPM>, and there are more instances containing “his <BPM>" than “her <BPM>".
Discussion: By evaluating different distributions of possessed BPM types, we can begin building an understanding of whose body parts are most often referred to in conversation. Our results indicate that individuals are more likely to discuss their own bodily experiences in online discourse than that of others. The higher frequencies of “his <BPM>" over “her <BPM>" instances are also interesting, considering well-established theories that women’s bodies are more heavily discussed and scrutinized in popular media Bordo (2023)—our results indicate that in spite of this, in terms of absolute numbers, men’s bodies may be referred to more often in everyday speech.
@@
Results: We find that there are certain BP word types that appear very frequently in all corpora—with 12 “my <BPM>" word types being shared in the top 20 across all corpora. However, we also observe variation in frequencies across corpora, indicating that we describe our body in different ways across different online contexts. Across all corpora, my heart and my head are among the most frequently mentioned “my <BPM>" word types. These body parts are likely central sources for people’s basic understandings of their embodied experiences, which is reflected in the prevalence of common figurative expressions such as my heart is broken and my head hurts. We also find that the blog dataset has a much stronger representation of body parts that are strongly related with human senses, such as my eyes (10.23% vs. 1.40%), my ears (1.19% vs. <0.1%), and my hands (3.32% <0.1%). This suggests that the personal narratives in blogs may be more focused on sensory, everyday experiences. Additionally, my hair and my face appear much more frequently in the TUSC tweet datasets than in the Spinn3r blog datasets, likely a result of personal grooming and appearance being more prevalent in social media updates. This rich divergence between common “my <BPM>" word types implies that users refer to their body differently when expressing themselves in different online mediums. (Table 5 in the Appendix shows the top 20 “my <BPM>" types in each corpus.)
@@
B4. Does the time of day/week/year impact whether we refer to our body?
Method: Each sample in the TUSCctry dataset has the exact timestamp at which it was posted. We use this data to examine whether “my <BPM>" usage is higher or lower at different times.
Results: Frequency of instances containing “my <BPM>" by day of week and month are shown in Figures 1 and 2. We find that “my <BPM>" instances peak during the summer and spring, decline steadily during the fall, and then stay relatively low during the winter. We also find that "<BPM>" instance frequency is highly dependent on different days of the week, rising steadily from Sunday to Wednesday and then declining from Wednesday to Saturday.
Discussion: The seasonal differences in “my <BPM>" usage in warmer months indicates that factors such as temperature, sunlight, and time spent outside could affect awareness and expression of one’s bodily experiences. The weekly rise of referral to one’s own body parts may reflect a renewed engagement with structured activities as the work week begins, while the decline could indicate fatigue or decreased energy as the week progresses, making it difficult to cultivate bodily awareness, consistent with documented patterns of weekly fatigue cycles in organizational research Zijlstra and Rook (2008). These results indicate that embodied language use is not static, but responds to environmental and social rhythms.
@@
Method: We take advantage of the geotagged meta data available for the TUSC tweets to evaluate the regional proportion of “my <BPM>" tweets.
Results: Figure 3 shows BPM use by city. We find that “my <BPM>" instances are more frequent in American tweets than Canadian tweets (TUSCctry dataset), and that “my <BPM>" instances are also used more in central cities than coastal cities (TUSCcity dataset).

### v2

@@ Abstract
This paper is the first investigation of the connection between emotion, embodiment, and everyday language in a large sample of natural language data. First, we created corpora of body part mentions (BPMs) in online English text (blog posts and tweets). These include a subset featuring human annotations for the emotions of the person whose body part is mentioned in the text. Next, we show that BPMs are common in personal narratives and tweets ($\sim$5% to 10% of posts include BPMs) and that their usage patterns vary markedly by time and location. Using word–emotion association lexicons and our annotated data, we show that text containing BPMs tends to be more emotionally charged than text without any BPMs. Finally, we show a strong and statistically significant correlation between body-related language and a variety of negative health outcomes. In sum, we argue that investigating the role of body-part related words in language can open up valuable avenues of future research at the intersection of NLP, the affective sciences, and the study of human wellbeing.
@@ 1 Introduction
Embodied cognition—the theory that human cognition is rooted in bodily experiences—has gained significant traction across cognitive science, psychology, linguistics, and philosophy. This framework suggests that our bodily experiences shape not only how we interact with our physical environment, but also how we process, represent, and share abstract concepts. A wide range of disciplines have demonstrated that the intersection of cognition and language is deeply intertwined with sensorimotor experiences. For example, a growing pool of research suggests that: individuals learn new concepts better when they can use their bodies to simulate the concepts Cook et al. (2008); Johnson-Glenberg et al. (2016); language processing involves mental simulation of physical actions Pulvermüller (2005); Glenberg and Kaschak (2002); and emotions emerge from interpretations of physiological signals through a process known as interoception Craig (2002); Barrett (2017).
In this paper, we ask the question: to what extent are our embodied experiences encoded in, and reflected through, everyday language? To this end, we use the concept of Body Part Mentions (BPMs), which we define as instances of language where words referring to parts of the body are used. [^1] Some linguistics paper use the term somatic reference for such mentions, and the term somatic unit or somatic phraseology for the vocabulary; however, somatic reference and somatic expression can be used more generally to express even non-language signals associated with the body. Even though BPMs are a relatively simple method for detecting embodiment-related language, we propose them as a useful initial tool for investigating the relationship between language and our embodied selves in everyday language. As such, we use BPMs in this paper for an exploratory investigation of how words which are semantically related to the human body can provide interesting insights in natural language use.
@@
We introduce two novel BPM corpora, Spinn3rBPM (a corpus of blog posts) and TUSCBPM (a corpus of tweets), and conduct a variety of experiments to show how they can be used to address three areas of inquiry on the connection between embodiment and natural language. First, we investigate the prevalence of body-part related words in everyday language, as well as how frequency of these words differs across different factors such as medium, gendered pronoun usage, time, and place. Second, we look at the relationship between BPMs and affect, motivated by the growing pool of research which posits that emotions originate from interpretations of physiological signals. Lastly, we propose and test the hypothesis that the degree of prevalence of BPMs in social media is indicative of aggregate-level health outcomes.
A better understanding of embodiment through language could support work in a range of research areas, especially NLP in health domains (where BPM-heavy corpora are frequent Chaturvedi et al. (2023)). the growing wave of interest in integrating embodiment within computational models of language (especially text-based LLMs, which rely on BPMs to access information about human embodiment, and struggle with many benchmarks of human cognition due to their lack of embodiment Chemero (2023)).
All of our code and data is available at our project repository. [^2] https://www.github.com/sohpei/bpms/
@@ 2.1 Embodiment, Affect, and Health
According to the theory of constructed emotion Barrett (2017), emotions emerge as interpretations of our bodies’ physiological signals. A key supporting argument for this theory is the wide range of research establishing a strong connection between interoception—the ability to feel internal bodily sensations Craig (2002)—and emotional wellbeing. Better interoceptive awareness has been shown to positively correlate with better emotional regulation Zamariola et al. (2019), emotional decision-making Dunn et al. (2010), and emotional granularity (the ability to distinguish different emotions) Feldman et al. (2024). These results indicate that awareness of our bodily experiences is crucial to our emotional welfare and dysfunctional interoception is a contributor to a variety of mental health conditions Khalsa et al. (2018).
Embodied experiences manifest frequently in everyday language. For example, descriptions of bodily experiences are frequently found in storytelling contexts Gallese and Wojciehowski (2011). But even outside of their role as explicit physical referents, BPMs may reveal a deeper connection to embodied phenomena. The theory of conceptual metaphor suggests that metaphors are a fundamental cognitive process Lakoff and Johnson (2008). According to this framework, metaphors help us understand abstract concepts by mapping them onto concrete, physical experiences. Treated this way, the widespread usage of body part words in language—even those applied in abstract contexts—are important for understanding how natural language is connected to the physical world from which it originates. Scholars in affective theory have argued that this is why expressions of intensely emotional experiences often use body parts and actions as a metaphor (e.g., my heart is broken, weight lifted off my shoulders); they are effective ways of grounding subjective experience in a shared embodied reality Kövecses (2003).
@@ 2.2 Embodiment and NLP
Language plays a key role in understanding the relationship between embodiment and affect. Emotional granularity is measured from the ability to identify and distinguish different emotions using words Tan et al. (2022), and interoception is often measured through an individual’s ability to describe their internal state Desmedt et al. (2022). A range of NLP projects have taken an interest in body part words for specific applications, such as identifying gendered representations in literature Silva et al. (2024), mapping bodily sensations for healthcare applications Wang et al. (2019), or building computational methods for detecting body parts involved in emotional processes to improve machine emotion recognition Zhuang et al. (2024). We note that other related work in NLP largely uses body-related language to answer adjacent research questions in specialized settings, rather than investigate the significance of BPMs themselves. Since this is the first-ever work investigating the significance of BPMs in everyday language using NLP methods, we aim to investigate whether the general relationship between affect and body parts suggested and observed in laboratory environments (i.e., the relationship between affect and body part words suggested by the theory of conceptual metaphor, or the relationship between affect and described bodily experiences suggested by the theory of constructed emotion), can be corroborated using textual corpora.
Many word–emotion association lexicons have been created: mostly for English: e.g., Bradley and Lang (1999), the NRC Emotion Lexicon Mohammad and Turney (2013, 2010), Warriner et al. (2013), The NRC VAD lexicon Mohammad (2018a); but also for some other languages: e.g., Moors et al. (2013) for Dutch, Võ et al. (2009) for German, and Redondo et al. (2007) for Spanish. The NRC Emotion Lexicon includes entries for whether a words is associated with eight categorical emotions: joy, sadness, fear, anticipation, anger, trust, disgust, and surprise (the Plutchik (1980) set). It includes entries for about 14,000 words. [^3] http://saifmohammad.com/WebPages/NRC-Emotion-Lexicon.htm The NRC VAD lexicon Mohammad (2018a) has valence, arousal, and dominance associations for about 20,000 English words. [^4] http://saifmohammad.com/WebPages/nrc-vad.html (Version 2 of the NRC VAD Lexicon was released recently, and it includes entries for about 44k words and 10k multiword expressions.)
In this work, we primarily use lexicons of word–emotion associations—which can be used to create accurate emotion arcs in streams of text Teodorescu and Mohammad (2023) or effectively distinguish emotional granularity Vishnubhotla et al. (2024) —as a computationally inexpensive and interpretable method for beginning to look at the relationship between affect, language, and embodiment.
@@ 3 BPM Corpora
For this work, we investigate two mediums for online, everyday language: blog posts and tweets. We consider each sentence from a blog post and each individual tweet from the tweet corpora as an instance. We created BPM corpora by extracting all instances that included at least one word referring to a body part from the Spinn3r personal blog datasets Burton et al. (2009, 2011), and the two TUSC datasets: TUSCctry (where tweets are geo-located to either the United States of America or Canada) and TUSCcity (where tweets are geo-located to cities in North America) Vishnubhotla and Mohammad (2022). This results in three final BPM corpora: Spinn3rBPM, TUSCctry-BPM, and TUSCcity-BPM. We will refer to the corpora made up of the rest of the instances as Spinn3rnoBPM and TUSCnoBPM.
We compile a list of body part words by including all the terms in the list used by Zhuang et al. (2024), which extracts BPM samples from the Spinn3r corpus to annotate samples for the presence of explicitly embodied emotion. [^5] https://www.collinsdictionary.com/us/word-lists/body-parts-of-the-body [^6] https: //www.enchantedlearning.com/wordlist/body.shtmlAdditionally, we include plural forms of terms (e.g., hearts, hands, eyes, etc.). We refer to the body part word forms as BP word types. The list of 295 BP word types we used is available on the project webpage.
Two issues need to be addressed when working with BPM corpora to study embodiment. First, some BP word types are ambiguous (some BPM instances may not actually be referring to a body part: e.g, ‘I will be back’). Second, it is useful to distinguish between the speaker referring to their own body vs. the speaker referring to someone else’s body. We are especially interested in mentions of one’s own body as a possibly useful indicator for well-being. While we are interested in the insights that all BPMs, including those with a more abstract/metaphorical connections to the human body, can offer on the relationship between embodiment and language, it would also be useful to distinguish instances where the speaker refers directly to their own body parts. and so we use a simple solution that effectively addresses both of the issues raised above. We created three subsets for each BPM corpus that only include the BPM instances preceded by possessive pronouns ‘my’, ‘your’, and ‘his/her/their’, respectively, and call these instances: possessed BPMs This helps us create separate corpora for first-, second-, and third-person references to body parts, and also excludes a vast majority of mentions that are not explicit references to body parts. For example, ‘I will be back’ would not be included in any of these corpora, but ‘my back hurts’ will be. We find that this approach delivers a high number of references to an individual’s body part (92% of 100 manually inspected instances). While this approach sacrifices recall by excluding possible references to body parts without possessive pronouns, we benefit from higher precision. [^7] For our experiments, we do not need all BPM instances, but rather just a large sample. Additionally, we also conduct some experiments with all BPM instances (the higher-recall and lower-precision corpus).
@@ 3.1 Emotion-Annotated BPM Corpus
We also leverage a specialized subset of Spinn3r, previously released by Zhuang et al. (2024). This subset contains sentences with BPMs annotated for the presence of explicitly embodied emotion, which is defined as "the physical experience of an emotion via our body" (i.e., "Julie pouted and rolled her eyes" is annotated as containing embodied emotion, but "Frank breathed heavily through his mouth after his run" is not). We refer to this subset in the rest of this paper as Spinn3rBPM-Zhuang. We extend this work by creating the first human-annotated dataset that explicitly identifies BPM ownership (whether the BPM refers to the speaker’s body or not) and the emotion of the BPM owner (joy, fear, etc.) as inferred by a human reader.
For full details of all datasets and data collection methods see Appendix˜A. Further details for the emotion annotation process, which includes a set of quality checks and aggregation methods for final emotion scores, can be found in Appendix˜B.
@@ 4 Research Questions About BPMs
Despite substantial evidence from medical research and psychology that points to connections between the mind and the rest of the body, as well as the connection between interoception and emotional well-being, there is little quantitative work using language to explore this connection. In this section and the next, we make use of large amounts of social media data, massive word–emotion lexicons, and the emotion-annotated corpus described in the previous section, to examine questions on how, when, and in what context we refer to our body parts in text (this section) and whether different body parts mentions tend to be used in different emotional contexts (next section). Since the questions in this section are relevant to the Body, we will index them as B1, B2, etc. The questions in the next section are related to Body and Affect, so we will index them as BA1, BA2, etc.
@@
Method: We did not find any past research on how common BPMs are in language. We do not even have a sense of the magnitude: do they occur in 0.01%, 0.0001%, 1%, 10%, 60% of utterances, or something else? To examine the extent to which body-related vocabulary is used in online language, we calculated the number of instances containing at least one BPM.
Results: See Table˜1. We find that a substantial proportion of instances contain at least one BPM: 10.4% of blog post sentences in the Spinn3r dataset, 6.4% and 7.3% of tweets in in TUSCctry and TUSCcity, respectively.
Discussion: The consistently high proportion of BPM-containing instances across our corpora emphasizes the ubiquity of body part references in online English text. The markedly larger proportion of samples containing BPMs in blog sentences than tweets implies that usage of BPMs may be more pervasive in personal narratives and longer-form text than in tweets.
@@
Method: Although BPMs can be used to study the general significance of body part words in language, we are also interested in examining the extent to which these body parts can be attributed to a particular person (e.g., her heart) vs. body part words that are not attributed to a human possessor (e.g., the heart of the matter). To do this, we introduce the concept of possessed BPMs, which can be attributed to a particular body using a possessive pronoun. We look at three general categories of possessed BPMs: first person instances including “my <BPM>”, second person instances including “your <BPM>”, and third person instances including “his/her/their <BPM>”. In each corpus, we determine the frequency of instances containing at least one instance of a possessed BPM.
Results: Table˜2 shows the results. In each corpus, the most common possessed BPM is “my <BPM>”, and there are more instances containing “his <BPM>” than “her <BPM”.
Discussion: By evaluating different distributions of possessed BPM types, we can begin building an understanding of whose body parts are most often referred to in conversation. Our results indicate that individuals are more likely to discuss their own bodily experiences in online discourse than that of others. The higher frequencies of “his <BPM>" over “her <BPM>" instances are also interesting, considering well-established theories that women’s bodies are more heavily discussed and scrutinized in popular media Bordo (2023)—our results indicate that in spite of this, men’s bodies may be referred to more often in everyday speech.
@@
Results: We find that there are certain BP word types that appear very frequently in all corpora—with twelve “my <BPM>” word types being shared in the top twenty across all corpora. However, we also observe variation in frequencies across corpora, indicating that we describe our body in different ways across different online contexts. Across all corpora, my heart and my head are among the most frequently mentioned “my <BPM>” word types. These body parts are likely central sources for people’s basic understandings of their embodied experiences, which is reflected in the prevalence of common figurative expressions such as my heart is broken and my head hurts. We also find that the blog dataset has a much stronger representation of body parts that are strongly related with human senses, such as my eyes (10.23% vs. 1.40%), my ears (1.19% vs. <0.1%), and my hands (3.32% <0.1%). This suggests that the personal narratives in blogs may be more focused on sensory, everyday experiences. Additionally, my hair and my face appear much more frequently in the TUSC tweet datasets than in the Spinn3r blog datasets, likely a result of personal grooming and appearance being more prevalent in social media updates. This rich divergence between common “my <BPM>” word types implies that users refer to their body differently when expressing themselves in different online mediums. (Table˜5 in the Appendix shows the top 20 “my <BPM>” types in each corpus.)
@@
B4. Does the time of the week/year impact the extent to which we refer to our body?
Method: Each sample in the TUSCctry dataset has a timestamp indicating the exact time at which it was posted. We use this data to examine whether “my <BPM>” usage is higher or lower at different times.
Results: Frequency of instances containing “my <BPM>” by day of week and month are shown in Figures 1 and 2. We find that “my <BPM>” instances peak during the summer and spring, decline steadily during the fall, and then stay relatively low during the winter. We also find that "<BPM>" instance frequency is highly dependent on different days of the week, rising steadily from Sunday to Wednesday and then declining from Wednesday to Saturday. We find a statistically significant decline in BPM usage from April onwards in the year and from Monday onwards throughout the week. More details on statistical tests can be found in Appendix˜H.
Discussion: The seasonal differences in “my <BPM>” usage in warmer months indicates that factors such as temperature, sunlight, and time spent outside could affect awareness and expression of one’s bodily experiences. The weekly rise of referral to one’s own body parts may reflect a renewed engagement with structured activities as the work week begins, while the decline could indicate fatigue or decreased energy as the week progresses, making it difficult to cultivate bodily awareness, consistent with documented patterns of weekly fatigue cycles in organizational research Zijlstra and Rook (2008). These results indicate that embodied language use is not static, but responds to environmental and social rhythms.
@@
Method: We take advantage of the geotagged meta data available for the TUSC tweets to evaluate the regional proportion of “my <BPM>” tweets.
Results: Figure 3 shows BPM use by city. We find that “my <BPM>” instances are used more in central cities than coastal cities. We also find that “my <BPM>” instances are more frequent in American tweets than Canadian tweets in the TUSCctry dataset, as shown in Figure˜4 (TUSCcity dataset). We find also find that the city of a post has a statistically significant effect on the usage of BPMs in a city. We also find that country has a slightly smaller but still statistically significant effect on usage, but that this pattern also persists across the different months represented in our dataset. Details on statistical tests, as well as a figure showing that usage difference between USA and Canada is consistent across months, can be found in Appendix˜D.